OpenStudy (ichbinpotato):

I need a step by step explanation for -5+i/2i Please help!

OpenStudy (welshfella):

are you sure you got that right?

OpenStudy (ichbinpotato):

Yes its supposed to be -5+i over 2i

OpenStudy (welshfella):

ok so we have an imaginary number as the denominator We need to convert it to a real number by a multiplication . Also we need to multiply the numerator by this same multiplier. Does that give you any ideas?

OpenStudy (ichbinpotato):

I'm honestly terrible at this. But, I'm guessing that it'd be multiplied by i (Which is -1) I'm seriously not sure of anything beyond that.

OpenStudy (welshfella):

yes actually i = square root of -1 so i^2 = -1 this willmake the denominator = -2 now multiply the top by i

OpenStudy (welshfella):

2 i * i = 2 i^2 = 2*-1 = -2

OpenStudy (ichbinpotato):

Oh, well now I feel like a dork. Is this all I had to do?

OpenStudy (ichbinpotato):

I'm assuming not since the top is still left, right?

OpenStudy (welshfella):

we have -5 + i (-5 + i)i ----- = --------- 2i -2

OpenStudy (welshfella):

Now you distribute the i over the (-5 + i)

OpenStudy (ichbinpotato):

So, -5i +i^2?

OpenStudy (welshfella):

right and i^2 = -1

OpenStudy (ichbinpotato):

Oh okay so -5i-1

OpenStudy (welshfella):

-1 - 5i ----- -2 Now we can divide the top and bottom by -1 to get rid of the negatives this gives us (1 + 5i ) / 2

OpenStudy (ichbinpotato):

Sorry, I'm trying to write all this down

OpenStudy (welshfella):

A complex number is usually written as a + bi The imaginary part comes second.

OpenStudy (welshfella):

no problem.

OpenStudy (ichbinpotato):

Mkay, so do we have to put (1+5i)/2 into that equation? (a+bi)

OpenStudy (welshfella):

Not really though i dont like fractions too much so i would write it as 0.5 + 2.5i which is in the form a + bi

OpenStudy (ichbinpotato):

Ah, I get it okay. I don't like fractions either which made this problem even more of a headache.

OpenStudy (ichbinpotato):

So would that be the end product?

OpenStudy (welshfella):

yes with these type of problems the way to go is to convert the denominator to a real number - remembering to multiply the numerator by the same multiplier in order to retain the same value

OpenStudy (welshfella):

If the denominator is a complex number you multiply by the Conjugate for example 1 + i ---- 2 + 3i - multiply by the conjugate of 2 + 3i which is 2 - 3i.

OpenStudy (welshfella):

because (2 + 3i)(2 - 3i) = 2^2 + 6i - 6i - 9 i^2 - note that the terms in i cancel out and the value becomes 4 - 9(-1) = 4 + 9 = 13 A real number

OpenStudy (ichbinpotato):

Thank you so much for such a detailed explanation! Seriously you've been such a great help!

OpenStudy (welshfella):

yw That's what this site is about.

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)